OpenStudy (anonymous):

Solve and graph inequality? 6.7 > =0.2x + 4.5 _

OpenStudy (john_es):

\[6.7-4.5\geq0.2x\Rightarrow2.2\geq 0.2x\Rightarrow11\geq x\] Without another variable, the graph is the straight line, that has enclosed all real numbers lower than or equal to 11.

OpenStudy (anonymous):

Um, could you try to explain more? Still a little confused. :|

OpenStudy (anonymous):

Opps, I forgot to put the negative sing infront of 0.2, so sorry, wouldn't that just mae the answer -11 though?

OpenStudy (john_es):

Well, the idea is solve the problem as if the equation was an equality. Group number in one side and variables in other. However, in inequalitys you must take care of the sign - with products and quotients. Firts, group numbers in the left side, \[6.7-4.5\geq-0.2x\] \[2.2\geq-0.2x\] Now divide both members by -0.2. When you divide or multiply by a number with the minus sign, the symbol of the inequality must change, \[\frac{2.2}{-0.2}\leq\frac{-0.2x}{-0.2}\] \[-11\leq x\] Then you can draw a line than begins in -11 and continues to the positive real numbers.

OpenStudy (anonymous):

Oh, okay, I understand now, thank you very much!

OpenStudy (john_es):

you're welcome ;)